A gas sample occupies 3.25 liters at 24.5 degrees Celsius and 1825 mm hg. Determine the temperature at which the gas will occupy 4250 ml at 1.50 atm
1
Expert's answer
2012-02-09T08:07:10-0500
V1=3,25 L V2=4250 ml=4,25
L
t1=24,5 C (273 +24,5)=297,5 K P2=1,5 atm =151,95 kPa
P=1825
mm hg =243,25 kPa t2=
???

P1V1=nRT1
P2V2=nRT2
nR=const

P1*V1/T1=P2*V2/T2

T2=P2*V2*T1/P1*V1

T2=(151,95*4,25*297,5)/243,25*3,23=242,8

t2=242,8